OpenStudy (anonymous):

The midpoint of UV is (5, –11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V

hartnn (hartnn):

co-ordinate of mid-point is average of co-ordinates of endpoints. let V=(x,y) so (3+x)/2 = 5 can u find x from here ?

OpenStudy (anonymous):

no..

hartnn (hartnn):

multiply 2 to both sides, what u get ?

OpenStudy (anonymous):

im confused..

hartnn (hartnn):

what is 5*2 =?

OpenStudy (anonymous):

10 lol.

hartnn (hartnn):

so 3+x=10 x=10-3 = ??

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

7?

hartnn (hartnn):

x=7. right :) now y. (5+y)/2=-11 y+5=-22 y= -22-5 = ??

OpenStudy (anonymous):

-27..

hartnn (hartnn):

thats your y. so your point V will be (7,-27)

OpenStudy (anonymous):

ooooohhhhhh thanksss haha

hartnn (hartnn):

welcome :)